# Slaves to the Law of Large Numbers: An Asymptotic Equipartition Property for Perplexity in Generative Language Models

### v3

@@ 3.4 Typical Sets are Vanishingly Small
The size of the typical set $\mathcal{P}_{M}^{n}$ is an exponentially vanishing fraction $02^{-\gamma n},~{}\gamma>$ 0 of all grammatically correct sequences.
The argument below closely tracks standard information-theoretic results (Cover & Thomas, 2005a) about the size of the typical set. We start by restating Proposition B.1 informally as:
$\Pr({\mathbf{Y}}_{n})$ $\equiv 2^{-nl_{M}({\mathbf{Y}}_{n})}\rightarrow 2^{-nh_{M}({\mathbf{Y}}_{n})}$ (20)
Let $T_{g}(n)$ be the total number of grammatically correct strings ${\mathbf{Y}}_{n}$ of length $n$. The average entropy $h_{M}({\mathbf{Y}})$ is a measure of the richness of the model that generated the string. Language models typically have a “temperature” parameter $T$ in the range $T\in[0,1]$ that can be used to tune the richness of the model’s output. Higher temperature means less “predictable” (i.e. less deterministic) and more “creative” (i.e. more random) outputs. This corresponds to higher values for $h_{M}({\mathbf{Y}})$.
Consider the maximally rich model of length $n$ that has the highest possible entropy $h_{M}({\mathbf{Y}}_{n})\equiv h_{max}(n)$ which corresponds to choosing all possible (grammatically correct) strings of length $n$ with equal probability [^3] Note that this does not mean that the token distributions $p_{m}(y)$ are uniform over all valid tokens at each step $m$; indeed tokens corresponding to richer sub-trees as shown in Fig. 1 should be chosen with higher probability to maximize the overall richness of the model’s output.. By definition, for the maximally rich model, the probability of choosing any valid string is $2^{-nh_{max}(n)}$ and this means the number of possible strings of length $n$ is:
$T_{g}(n)$ $\equiv 2^{nh_{max}(n)}$ (21)
We can take (21) as the defining relation for $h_{max}(n)$. In practical LLMs, even at the highest temperature setting i.e. $T=1$, the token probability is typically concentrated in a small number of tokens and the token probabilities are determined by far more complex rules than the richness of their sub-trees. Let us formalize this as an assumption.
Assumption 3.1.
Practical LLMs are not entropy-maximizing. Any language model M has an entropy gap $\Delta h>0$ such that the following holds for all sufficiently long strings:
$h_{M}({\mathbf{Y}}_{n})$ $<h_{max}(n)-\Delta h,~{}\forall{\mathbf{Y}}_{n}\in\mathcal{T}^{n}$ (22)
Proof of Proposition 3.3.
From (20), for $n\gg 1$, we have $\Pr({\mathbf{Y}}_{n})\equiv 2^{-nl_{M}({\mathbf{Y}}_{n})}\to 2^{-nh_{M}({\mathbf{Y}}_{n})}>2^{-n(h_{max}(n)-\Delta h)}$ where we used Assumption 3.1 in the last step. Since this holds for any typical string of length $n$, we must have for small $\epsilon>0$ and correspondingly large $n$:
$\left|\mathcal{P}_{M}^{n}(\epsilon)\right|$ $<2^{n\left(h_{max}(n)-\Delta h)\right)}\equiv 2^{-n\Delta h}\left|T_{g}(n)\right|$ (23)
Thus, the typical set $\mathcal{P}_{M}^{n}(\epsilon)$ represents an exponentially vanishing fraction of all grammatically correct strings. ∎
We emphasize that all we needed to establish Proposition 3.3 was Assumption 3.1 which only requires that the model assigns higher probability to some sentences than to others.
3.5 Over-Typical Sets are Vanishingly Smaller Still
@@
$\mathcal{V}_{M}^{n}(\epsilon)$$\doteq\left\{{\mathbf{Y}}_{n}:\left(h_{M}({\mathbf{Y}}_{n})-l_{M}({\mathbf{Y}}_{n})\right)>\epsilon\right\}$
$\equiv\left\{{\mathbf{Y}}_{n}:\Pr({\mathbf{Y}}_{n})>2^{-n\left(h_{M}({\mathbf{Y}}_{n})-\epsilon\right)}\right\}$ (24)
@@
The size of the over-typical set $\mathcal{V}_{M}^{n}(\epsilon)$ is an exponentially vanishing fraction of all grammatically correct sequences.
@@
For any ${\mathbf{Y}}_{n}\in\mathcal{V}_{M}^{n}(\epsilon)$, we have from (24) and (22):
$\Pr({\mathbf{Y}}_{n})$ $>2^{-nh_{max}(n)}\times 2^{n(\Delta h+\epsilon)}$ (25)
Therefore, we must have
$\left|\mathcal{V}_{M}^{n}(\epsilon)\right|$ $<\left|T_{g}(n)\right|2^{-n(\Delta h+\epsilon)}$
where we used the definition of $h_{max}$ in (21). ∎
We have shown that almost all long grammatically correct strings are under-typical for any language model M. We illustrate these ideas for a simple Markov language model in Appendix A.
@@ 4 Experiments with Open-Source Models
We performed a series of experiments to verify the ideas described in Section 3. The basic idea is to use a local instance of a pre-trained open-source model to evaluate the empirical entropy of a given string using the model’s conditional probability distributions and compare the result with the log-perplexity of the string again calculated using the model’s probability distributions. We ran experiments on the smaller GPT-2 model (Radford et al., 2019) as well as the more recent, larger and more capable Llama 3.1 8B (Touvron et al., 2023) model.
@@ 4.1 Perplexity of Self-Generated Strings
We initialized the prompt string and the seed of random number generator with fixed values to allow for later reproduction. We then used top-k sampling with $k=100$ to generate the probability distribution for the next token. We normalized the probabilities of the top-k tokens to sum to unity. We then created and saved a table consisting of token ID, token string and selection probability of each of the top $k$ tokens. Finally, a random next token is chosen by sampling from the normalized distribution, and the new token is appended to the prompt to repeat the process for $N_{max}$ total tokens. We want the number of tokens $N_{max}$ to be large enough to observe the asymptotic behavior of the perplexity of the generated string. In practice, $N_{max}$ is constrained by certain limitations of the models as we discuss in the sequel.
The random token $Y_{n}$ chosen at each step along with the probability distribution $p_{n}(y)$ from which they were chosen were saved into a json file for later processing. We ran this experiment many times for different seed values for the random number generator, different choices of sampling distributions and a number of different initial prompts.
For each generated string, we can then analyze the information archived in the json file to test our claim in Proposition 3.1. Specifically, for each token $N\in 1\dots N_{max}$ we calculated the log-perplexity $l_{M}({\mathbf{Y}}_{N})$, empirical entropy $h_{M}({\mathbf{Y}}_{N})$ and the log-deviation $\lambda_{M}({\mathbf{Y}}_{N})$ of the substring ${\mathbf{Y}}_{N}$ consisting of the first $N$ tokens of the generated string.
@@ 4.2 Perplexity of Externally Generated Strings
Stating with $n=1$, we list the probability distribution $p_{n}(y)$ of the next token $Y_{n}$ for the model using top-k with $k=100$ sampling for the $n$’th token. We can then calculate the empirical entropy $h_{M}({\mathbf{Y}}_{N})$ and log-deviation $\lambda_{M}({\mathbf{Y}}_{N})$ as in Section 4.1. However, to calculate the log-perplexity $l_{M}({\mathbf{Y}}_{N})$, we also need the probability that our model will output the actual next token $Y_{n}$ from the string ${\mathbf{Y}}$ which may not be included in the list of top-$k$ tokens [^4] Note that our choice of top-k sampling is a compromise. The number of tokens in GPT-2 or Llama is quite large ($|\mathcal{T}|\approx 40,000$). If we exhaustively list the full probability distribution of all tokens, our experiment becomes very computationally expensive and slow.. When this happens, we replaced the lowest probability entry in the list of top-k tokens with the actual next token from our string along with its probability.
The normalized probability of each of the top-k tokens along with their token-ID and token strings, as well as the probability and ID of the actual next token are saved in a json file. This json file is then processed in exactly the same way as in Section 4.1 to compute the log-perplexity $l_{M}({\mathbf{Y}}_{N})$, empirical entropy $h_{M}({\mathbf{Y}}_{N})$ and the log-deviation $\lambda_{M}({\mathbf{Y}}_{N})$ of the sub-string ${\mathbf{Y}}_{N}$ for the GPT-2 model for $N=1\dots N_{max}$.
@@ 4.3 Results and Discussion
Figure 2 shows the log-perplexity and empirical entropy for sub-strings ${\mathbf{Y}}_{N}$ for the GPT-2 or Llama 3.1 8B model as a function of $N$ for several text strings generated by the same model. The plots also shows the range $h_{M}({\mathbf{Y}}_{N})\pm\lambda_{M}({\mathbf{Y}}_{N})$ and $h_{M}({\mathbf{Y}}_{N})\pm 2\lambda_{N}({\mathbf{Y}}_{N})$. We see that the log-perplexity $l_{M}({\mathbf{Y}}_{N})$ stays within the range $h_{M}({\mathbf{Y}}_{N})\pm 2\lambda_{M}({\mathbf{Y}}_{N})$ consistently in all cases. The asymptotic convergence is more clearly seen in the long strings from the Llama 3.1 8B model [^5] The GPT-2 model cannot generate similarly long texts. in Fig. 2(c), 2(d) . We find that very long synthetic texts such as those in Fig. 2(c), 2(d) can be ungrammatical, repetitive and generally of poor quality, as reported elsewhere (Holtzman et al., 2020). However, Proposition 3.1 holds regardless of the quality of the text as seen in Fig. 2.
@@
(b) Llama 3.1 8B, $N_{max}=1024$.
(c) Llama 3.1 8B, $N_{max}=10240$.
(d) Llama 3.1 8B, $N_{max}=10240$.
Figure 2: Log-perplexity and empirical entropy for GPT-2 or Llama 3.1 8B of self-generated text, top-100 sampling.
Figure 3 shows the log-perplexity and empirical entropy for the Llama 3.1 8B model for two long, non-Llama generated strings. The text used in Fig. 3(a) is an excerpt from a recent article (Pelly, 2025) in Harper’s Magazine, a respected literary publication. The string used in Fig. 3(b) was generated by ChatGPT o1 Pro (OpenAI, 2024).
@@
We see from Fig. 3(a) and Fig. 3(b) that the log-perplexity $l_{M}({\mathbf{Y}}_{N})$ for Llama 3.1 8B is several standard deviations $\lambda_{M}({\mathbf{Y}}_{N})$ larger than the empirical entropy $h_{M}({\mathbf{Y}}_{N})$ showing that the corresponding strings are clearly under-typical for the Llama 3.1 8B model. While this non-convergence is not required by our theory - we have not considered whether the converse to Proposition 3.1 is true - it is strongly suggested by Propositions 3.3 and 3.4 if we assume that the strings in Fig. 3 are randomly sampled from the set of all grammatically correct texts.
@@ 4.4 More Catholic than the Pope
This brings us to the remarkable plots in Fig. 4 which shows the log-perplexity of excerpts from classic texts for Llama 3.1 8B, specifically from Alice in Wonderland and The Great Gatsby in Figs. 4(a) and 4(b) respectively. The average entropy for the token distributions in Fig. 4(a) is an astonishingly small $0.11$ bits, which means that the model on average assigns a very high probability ($\sim 99\%$) to the precise sequence of tokens from the excerpt for its predicted next tokens. In other words, this excerpt has been memorized by the Llama 3.1 8B model. The same observation also applies to Fig. 4(b). We also see from both plots that the log-perplexity is several standard deviations below the empirical entropy making the excerpts slightly, but unambiguously, over-typical for the Llama-8B model: the model judges these texts to be more perfect than its own creations!
From Proposition 3.4, we can conclude that these excerpts are among the small number of very special strings to be so memorized by the model. Clearly, these texts were part of the training set of the Llama-8B model and they were esteemed highly by its cost function.
Interestingly, the over-typicality of these excerpts means that it is pretty much impossible for the Llama-8B model itself to generate them verbatim - even with a $99\%$ probability that each next token will exactly match the excerpt, a randomly generated string can be expected to diverge from the excerpt after only about $100$ tokens.
Additional plots and more details about these experiments are presented in Appendix C.1.
5 Practical Applications
Consider the problem of AI text detection i.e. determining whether a string ${\mathbf{Y}}$ was generated by a model M. The intuition behind (9) suggests a simple method for doing this: check if the perplexity is small enough. Specifically, declare ‘yes’ if $l_{M}({\mathbf{Y}})<l_{th}$ for a threshold $l_{th}$. The same test can also be used for the problem of LLM dataset inference i.e. determining whether a string ${\mathbf{Y}}$ was used as part of the training dataset for the model M.
While, this idea is not new and indeed variations of it are behind popular solutions for both problems (Mitchell et al., 2023; Maini et al., 2024), our results in this paper provide a rigorous theoretical foundation for it. It also provides a principled way to choose the classification threshold $l_{th}$ which previously had to be computed empirically (Carlini et al., 2021). As an example, with $l_{th}=h_{M}({\mathbf{Y}})+3\lambda_{M}({\mathbf{Y}})$, Fig. 2 shows that we can accurately distinguish between synthetic text generated by Llama 3.1 8B and non-Llama text such as in Fig. 3 even with relatively short strings of $1000$ tokens. We can also provide performance guarantees. Indeed, a simple lower-bound on the probability of a false negative (i.e. failing to detect a Llama-generated string) can be obtained from (18). Similar comments apply to the dataset inference problem: while it is well-known that text strings memorized by an LLM have low perplexity, our theory allows us to rigorously derive threshold values and performance guarantees. Extrapolating from Fig. 4, we posit that if a long excerpt from a human-written book is over-typical for a language model, it is a near-statistical certainty that the book was part of the training data set for that model.

### v4

@@ 4.2 The Essential Smallness of Typical Sets
Typical sets are small. Under Assumption 4.1, the size of the pruned typical set $\mathcal{P}_{M,G}^{n}$ is an exponentially vanishing fraction of all grammatically correct sequences $\mathcal{G}(n)$.
Proof.
From Proposition 3.1, for sufficiently large $n$, we have $|l_{M}({\mathbf{Y}}_{n})-h_{M}({\mathbf{Y}}_{n})|\leq\frac{\Delta g}{2}$. Using this observation for ${\mathbf{Y}}_{n}\in\mathcal{E}_{2}(n)$, we have $l_{M}({\mathbf{Y}}_{n})\leq g(n)-\frac{\Delta g}{2}$. Then we have:
$1$ $\geq\Pr(\mathcal{E}_{2}(n))\equiv\sum_{{\mathbf{Y}}_{n}\in\mathcal{E}_{2}(n)}2^{-nl_{M}({\mathbf{Y}}_{n})}\geq\sum_{{\mathbf{Y}}_{n}\in\mathcal{E}_{2}(n)}2^{-n\left(g(n)-\frac{\Delta g}{2}\right)}\equiv\frac{|\mathcal{E}_{2}(n)|}{|\mathcal{G}(n)|}2^{\frac{n\Delta g}{2}}\geq\frac{\rho|\mathcal{P}_{M,G}^{n}(\epsilon)|}{|\mathcal{G}(n)|}2^{\frac{n\Delta g}{2}}$ (25)
Thus, we have shown $|\mathcal{P}_{M,G}^{n}(\epsilon)|\leq|\mathcal{G}(n)|2^{-n\left(\frac{\Delta g}{2}-\frac{\log_{2}(\rho)}{n}\right)}$. ∎
@@
$\mathcal{V}_{M}^{n}(\epsilon)$ $\doteq\left\{{\mathbf{Y}}_{n}\in\mathcal{G}(n):l_{M}({\mathbf{Y}}_{n})\leq h_{M}({\mathbf{Y}}_{n})-\epsilon<g(n)-\epsilon\right\}$ (26)
@@
Over-Typical sets are smaller still. The size of the over-typical set $\mathcal{V}_{M}^{n}(\epsilon)$ is an exponentially vanishing fraction of all grammatically correct sequences.
@@
The argument is similar to the proof of Proposition 4.2:
$1$ $\geq\Pr\left({\mathbf{Y}}_{n}\in\mathcal{V}_{M}^{n}(\epsilon)\right)\equiv\sum_{{\mathbf{Y}}_{n}\in\mathcal{V}_{M}^{n}(\epsilon)}2^{-nl_{M}({\mathbf{Y}}_{n})}\geq\left|\mathcal{V}_{M}^{n}(\epsilon)\right|2^{-n(g(n)-\epsilon)}\equiv\frac{\left|\mathcal{V}_{M}^{n}(\epsilon)\right|}{\left|\mathcal{G}(n)\right|}2^{n\epsilon}$ (27)
Thus we have $\left|\mathcal{V}_{M}^{n}(\epsilon)\right|\leq\left|\mathcal{G}(n)\right|2^{-n\epsilon}$ which proves the result. ∎
For Assumption 4.1 to be violated, we need the perplexities of almost all typical strings to be almost equal; conversely, any model that chooses some strings with higher probabilities than others must satisfy Assumption 4.1 and therefore Proposition 4.2. Together with Proposition 4.3, this means that almost all grammatical strings are undertypical for any practical LLM.
4.3 Constriction by Staying True to Type
We now present some example consequences of Proposition 4.2 as a payoff for the preceding analysis. Informally, Proposition 3.1 says that all models must stay “true to type”, and Proposition 4.2 says that this narrowly circumscribes the range of a model’s possible behaviors and outputs.
Consider the augmented language model $G=(M,D)$ shown in Fig. 2, where the output of a model $M$ is screened by a grammar-checking device $D$ which defines the dictionary $\mathcal{G}(n)$; if the model’s output $\mathbf{Y}$} is not grammatically correct, then it is prompted again to produce a new random string $\mathbf{Y}$} until it generates a grammatical string [^3] Note that we assume that the device $D$ is deterministic, which means our analysis does not allow the grammar checking device to be another stochastic LLM.. The construction in Fig. 2 implicitly assumes a decidable grammar, which means it describes a “recursive language” in Type 1 of the Chomsky hierarchy - the same category natural languages are believed to belong to (Chomsky, 2014).
Figure 2: Language Model Augmented with a Grammar Checking Device.
Since $p_{G}(n)>p_{G}>0$, the augmented model $G$ is guaranteed to generate a grammatical string in a finite time with probability one and the probability that model $G$ generates a grammatical string ${\mathbf{Y}}_{n}$ is $\Pr({\mathbf{Y}}_{n}|{\mathbf{Y}}_{n}\in\mathcal{G}(n))$. Furthermore, we see from (23) that the augmented model $G$ will only generate outputs from the set $\mathcal{P}_{M,G}^{n}(\epsilon)$ with high probability. We can then interpret the Proposition 4.2 as a restriction on what the augmented model $G$ can generate as illustrated in the following examples:
  1. Ex 1
    The Python programming language. We define $\mathcal{G}(n)$ as the set of all syntactically correct Python code that compiles (or interprets) without error. While we know LLMs are capable of generating complex Python code, Proposition 4.2 asserts that any model is limited to generating a vanishing fraction of all valid Python programs.
  2. Ex 2
    Proofs in First Order Logic (FOL). We define $\mathcal{G}(n)$ as the set of valid mathematical proofs in FOL. A model used to generate proofs in FOL is limited to generating a vanishing fraction of all correct proofs.
  3. Ex 3
    Lyrical Poetry. We define $\mathcal{G}(n)$ as the set of stories consisting of a sequence of haikus. A model is limited to a vanishing fraction of all such possible stories.
@@ 5 Experiments with Open-Source Models
We performed a series of experiments to verify the ideas described in Sections 3 and 4. The basic idea is to use a local instance of a pre-trained open-source model to evaluate the empirical entropy of a given string using the model’s conditional probability distributions and compare the result with the log-perplexity of the string again calculated using the model’s probability distributions. We used the smaller GPT-2 model (Radford et al., 2019) as well as the larger and more capable Llama 3.1 8B (Touvron et al., 2023) model.
@@ 5.1 Perplexity of Self-Generated Strings
We initialized the prompt string and the seed of random number generator with fixed values to allow for later reproduction. We then used top-k sampling with $k=100$ to generate the probability distribution for the next token. We normalized the probabilities of the top-k tokens to sum to unity. We then created and saved a table consisting of token ID, token string and selection probability of each of the top $k$ tokens. Finally, a random next token is chosen by sampling from the normalized distribution, and the new token is appended to the prompt to repeat the process for $N_{max}$ total tokens. We want the number of tokens $N_{max}$ to be large enough to observe the asymptotic behavior of the perplexity of the generated string.
The random token $Y_{n}$ chosen at each step along with the probability distribution $p_{n}(y)$ from which they were chosen were saved into a json file for later processing. We ran this experiment many times for different seed values for the random number generator, different choices of sampling distributions and initial prompts.
For each generated string, we can then test our claim in Proposition 3.1. Specifically, for each token $N\in 1\dots N_{max}$ we calculated the log-perplexity $l_{M}({\mathbf{Y}}_{N})$, empirical entropy $h_{M}({\mathbf{Y}}_{N})$ and the log-deviation $\lambda_{M}({\mathbf{Y}}_{N})$ of the substring ${\mathbf{Y}}_{N}$ consisting of the first $N$ tokens of the generated string.
@@ 5.2 Perplexity of Externally Generated Strings
Stating with $n=1$, we list the probability distribution $p_{n}(y)$ of the next token $Y_{n}$ for the model using top-k with $k=100$ sampling for the $n$’th token. We can then calculate the empirical entropy $h_{M}({\mathbf{Y}}_{N})$ and log-deviation $\lambda_{M}({\mathbf{Y}}_{N})$ as in Section 5.1. However, to calculate the log-perplexity $l_{M}({\mathbf{Y}}_{N})$, we also need the probability that our model will output the actual next token $Y_{n}$ from the string $\mathbf{Y}$} which may not be included in the list of top-$k$ tokens [^4] Our choice of top-k sampling is a compromise. The number of tokens in GPT-2 or Llama is quite large ($|\mathcal{T}|\approx 40,000$). If we exhaustively list the full probability distribution of all tokens, the experiment becomes very computationally expensive and slow.. When this happens, we replaced the lowest probability entry in the list of top-k tokens with the actual next token from our string along with its probability.
The normalized probability of each of the top-k tokens along with their token-ID and token strings, as well as the probability and ID of the actual next token are saved in a json file. This json file is then processed in exactly the same way as in Section 5.1 to compute $l_{M}({\mathbf{Y}}_{N}),~h_{M}({\mathbf{Y}}_{N})$ and $\lambda_{M}({\mathbf{Y}}_{N})$ of the sub-strings ${\mathbf{Y}}_{N},~N=1\dots N_{max}$ for the GPT-2 or Llama 3.1 8B model.
@@ 5.3 Results and Discussion
(b) Llama 3.1 8B, $N_{max}=10240$.
Figure 3: Log-perplexity & empirical entropy for GPT-2 or Llama 3.1, self-generated text, top-100 sampling.
Figure 3 shows the log-perplexity and empirical entropy for sub-strings ${\mathbf{Y}}_{N}$ for the GPT-2 or Llama 3.1 8B model as a function of $N$ for several text strings generated by the same model. The plots also shows the range $h_{M}({\mathbf{Y}}_{N})\pm\lambda_{M}({\mathbf{Y}}_{N})$ and $h_{M}({\mathbf{Y}}_{N})\pm 2\lambda_{N}({\mathbf{Y}}_{N})$. We see that the log-perplexity $l_{M}({\mathbf{Y}}_{N})$ stays within the range $h_{M}({\mathbf{Y}}_{N})\pm 2\lambda_{M}({\mathbf{Y}}_{N})$ consistently in both cases. The asymptotic convergence is more clearly seen in the long string from the Llama 3.1 8B model [^5] The GPT-2 model cannot generate similarly long texts. in Fig. 3(b). Additional plots showing the asymptotic behavior are presented in Appendix LABEL:sec:more_exp. However, we find that very long synthetic texts such as those in Fig. 3(b) can be ungrammatical and generally of poor quality, as reported elsewhere Holtzman et al. (2020).
While Proposition 3.1 holds regardless of the quality of the text, its practical relevance depends on whether the predicted asymptotic behavior is observable in normal-length texts. Figure 4 shows the log-perplexity $l_{M}({\mathbf{Y}}_{N})$ against the empirical entropy $h_{M}({\mathbf{Y}}_{N})$ for the Llama-3.1 8B model for several text strings generated by the same model. A total of $450$ strings were generated using top-$100$ sampling from $15$ distinct prompts and $30$ different random seeds for each prompt. Of these, abnormally short strings of less than $100$ tokens were excluded from the analysis. The remaining $383$ strings had an average length of $685$ tokens and a standard deviation of $321$. The clustering of the log-perplexity around the $45^{\circ}$ line in Fig. 4 provides direct visual affirmation for Proposition 3.1.
Figure 4: Log-perplexity & empirical entropy of “normal-length” text generated by Llama-3.1 8B.
The fact that the perplexity of the generated strings vary over a wide range in Fig. 4 also provides direct evidence for the Assumption 4.1 that the Llama-3.1 8B model is not entropy-maximizing and therefore must have vanishingly small typical sets per Proposition 4.2. Additional evidence for Propositions 4.2 and 4.3 is in Fig. 5 that analyses two long, non-Llama generated strings. The text used in Fig. 5(a) is an excerpt from a recent article Pelly (2025) in Harper’s Magazine, a respected literary publication. The text in Fig. 5(b) was generated by ChatGPT o1 Pro OpenAI (2024). We see that the log-perplexity $l_{M}({\mathbf{Y}}_{N})$ for Llama 3.1 8B is several standard deviations $\lambda_{M}({\mathbf{Y}}_{N})$ larger than the empirical entropy $h_{M}({\mathbf{Y}}_{N})$ for both strings in Fig. 5 showing that these strings are clearly under-typical for the Llama 3.1 8B model. If we assume that the strings in Fig. 5 are randomly sampled from the set of all grammatically correct texts, then this under-typicality is a simple consequence of the smallness of the Llama-3.1 8B model’s typical sets.
@@ 5.4 More Catholic than the Pope
This brings us to the remarkable plots in Fig. 6 which shows the log-perplexity of excerpts from classic texts for Llama 3.1 8B, specifically from Alice in Wonderland and The Great Gatsby in Figs. 6(a) and 6(b) respectively. The average entropy for the token distributions in Fig. 6(a) is an astonishingly small $0.11$ bits, which means that the model on average assigns a very high probability ($\sim 99\%$) to the precise sequence of tokens from the excerpt for its predicted next tokens. In other words, this excerpt has been memorized by the Llama 3.1 8B model in the sense of eidetic memorization Carlini et al. (2021): with zero temperature, it will output this text verbatim. The same observation also applies to Fig. 6(b).
We also see from both plots that the log-perplexity is several standard deviations below the empirical entropy making the excerpts slightly, but unambiguously, over-typical for the Llama-8B model: the model judges these texts to be more perfect than its own creations! From Proposition 4.3, we can conclude that these excerpts are among the small number of very special strings to be so memorized by the model. Clearly, these texts were part of the training set of the Llama-8B model and they were esteemed highly by its cost function.
@@
In summary, we see from Figs. 3 and 4 that synthetic texts generated by an LLM are typical for the model, while most external texts such as in Fig. 5 are undertypical. By definition, over-typical strings are very special such as the literary classics in Fig. 6 that have been memorized by the model.
6 Practical applications: the capybara problem
A simple and direct application of the ideas in this work is for testing whether a given piece of text has “abnormally” high or low perplexity for a language model. Such a test is relevant to many important practical problems. Low perplexity is commonly used as one marker of synthetic text in AI text detection tools (Wu et al., 2025) by testing $l_{M}({\mathbf{Y}}_{n})\begin{subarray}{c}<\\ >\end{subarray}l_{th}$ against a threshold $l_{th}$. However, it turns out that choosing a suitable classification threshold $l_{th}$ is difficult - the so-called “capybara problem” (Hans et al., 2024).
Our theory shows a principled way to do this: with $l_{th}=h_{M}({\mathbf{Y}}_{n})+3\lambda_{M}({\mathbf{Y}}_{n})$, Fig. 4 shows that we can accurately distinguish between synthetic text generated by Llama 3.1 8B and non-Llama text such as in Fig. 5 even with relatively short strings of $\approx 500$ tokens. The key insight is to choose variable rather than fixed thresholds, i.e. different $l_{th}$ for different strings ${\mathbf{Y}}_{n}$.
Our theory can also provide some performance guarantees: e.g. a simple upper-bound on the false negative rate (i.e. failing to detect a Llama-generated string) can be obtained from (17). Note, however, that we cannot provide bounds on false positive because the converse to Proposition 3.1 does not hold in general without additional assumptions.
Similar comments apply to the dataset inference problem (Maini et al., 2024): while it is well-known (Carlini et al., 2021) that text strings memorized by an LLM have low perplexity, our theory allows us to rigorously derive threshold values and performance guarantees. Extrapolating from Fig. 6, we posit that if a long excerpt from a human-written book is over-typical for a language model, it is a near-statistical certainty that the book was part of the training data set for that model.